# Standards-Compliant DM-RS Allocation via Temporal Channel Prediction for Massive MIMO Systems

### v1

@@ III-A Dataset Generation
We utilized the NVIDIA Sionna Ray Tracing to generate channel data [NVIDIA], as it more accurately reflects real-world propagation characteristics and physical environments than statistical channel models. Specifically, we employed the simple_street_canyon_with_cars scene included in the framework. As illustrated in Fig. 3, this scenario features a road between buildings, where vehicles move along both lanes. All vehicles move at a uniform speed $v\in\{60,65,\dots,80\}$ km/h. The BS is installed on one of the buildings and the UE is positioned in one of the vehicles. The BS has an 8×8 antenna array, while the UE has a 2×1 antenna array, forming a massive MIMO system. We consider an FDD scenario operating in NR band n1, as specified in Table 5.2-1 of [TS138101], using uplink and downlink carrier frequencies of 1.95 GHz and 2.14 GHz, respectively. The numerology is fixed to N=1, such that one subframe corresponds to a single slot. The maximum number of rays was set to 20, with a maximum ray depth of 4 and a bandwidth of 100 MHz.
For each velocity $v$, we obtained the uplink and downlink channel matrices for every symbol within 400 consecutive slots, each comprising 14 symbols and 64 subcarriers. Additionally, we augmented the original dataset by adding Gaussian noise with a magnitude of approximately 0.05 relative to the original channel matrices, in order to emulate practical channel estimation errors caused by uplink noise and other impairments. This augmentation generated an additional 4,000 slots. This resulted in a total dataset comprising 6,000 slots. Subsequently, for each slot, the optimal DM-RS allocation index was determined via exhaustive evaluation of all candidate configurations using Sionna link-level simulations, based on the objective function defined in (5). The dataset was then divided into training, validation, and test sets using an 8:1:1 ratio.

### v2

@@ III-A Dataset Generation
We utilized the ray-tracing module of NVIDIA Sionna to generate channel data [NVIDIA], as it more accurately reflects real-world propagation characteristics and physical environments than statistical channel models. We employed the simple_street_canyon_with_cars scene, where vehicles move along a two-lane road between buildings, as shown in Fig. 3. The BS is mounted on a building, and the UE is placed in a vehicle moving at $v\in\{60,65,\dots,80\}$ km/h. The BS has an 8×8 antenna array, while the UE has a 2×1 antenna array, forming a massive MIMO system. We consider an FDD scenario operating in NR band n1, as specified in Table 5.2-1 of [TS138101], using uplink and downlink carrier frequencies of 1.95 GHz and 2.14 GHz, respectively. The numerology is fixed to N=1, such that one subframe corresponds to a single slot. The maximum number of rays was set to 20, with a maximum ray depth of 4 and a bandwidth of 100 MHz.
To emulate practical estimation errors, we added Gaussian noise with magnitude $\approx$0.05 relative to the original channels, generating 4,000 additional slots and yielding 6,000 slots in total. For each slot, the optimal DM-RS allocation was obtained via exhaustive search using Sionna link-level simulations based on (5), and the dataset was split into training/validation/test sets with an 8:1:1 ratio.
IV Experimental Results
IV-A Considered Schemes
CPRS (ViViT): The proposed scheme described in Section III. It adopts a tubelet embedding with a patch size of $(1,8,8)$ and a Transformer backbone consisting of 8 encoder layers, each with 16 attention heads and an embedding dimension of 128. These architectural parameters were selected based on ablation studies, where this configuration consistently achieved the best performance among the considered ViViT variants.
CPRS (CNN): A CPRS variant based on a 3D convolutional neural network, where all $\mathbf{X}_{i}$ are concatenated into a four-channel input. The network consists of three Conv3D layers with 16, 32, and 64 filters, respectively, each followed by batch normalization and 3D max pooling, and a 128-unit dense layer at the output. This configuration was chosen as it yielded the highest performance among evaluated CNN-based architectures.
Disjoint (ViViT/CNN): In the disjoint baseline, the downlink CSI is first predicted using a ViViT- or CNN-based channel predictor [YANG2020]. To ensure a fair comparison, the ViViT predictor adopts the same patch size $(1,8,8)$ and backbone depth (8 layers with 16 attention heads) as CPRS (ViViT), while the CNN predictor uses the same 3D-CNN architecture as CPRS (CNN). The predicted CSI is then used to optimize the DM-RS pattern by maximizing channel estimation accuracy, following the procedure in [MASHHADI2021Prun] with adaptations for NR compliance.
DRL [KIM2023]: A CNN-based deep Q-network employing an $\epsilon$-greedy policy ($\epsilon=0.3$), where the uplink CSI serves as the state, 3GPP-compliant DM-RS patterns constitute the action space, and throughput is used as the reward.
Maximize Data Symbols: A heuristic scheme that maximizes the number of transmitted data symbols by minimizing the number of DM-RS symbols per slot ($n_{p}=1$).
Random Average: A scheme that selects all possible DM-RS allocations with equal probability. The reported performance corresponds to the average over all possible allocations.
Best: The ground-truth DM-RS allocation that achieves the optimal data throughput obtained through simulations.